OpenStudy (anonymous):

At 3:00pm, Catanya and Chad has already made 46 candles. By 5:00pm, the total reached100 candles. Assuming a constant production rate, at what time did they make their 82nd candle? please show work!

OpenStudy (gw2011):

I get an answer that they make their 82nd candle at 4:20pm. Here is how I got that answer: 100 - 46 = 54 additional candles were made in 2 hours 54/2 = 27 candles are made per hour since it is a constant production rate Then 1 hour later (4:00pm) 47 + 27 = 73 candles were made 82 - 73 = 9 more candles need to be made to get the 82nd candle Therefore, this takes 9/27 or 1/3 of an hour and (1/3)(60 minutes) = 20 mins. Therefore, 4:00pm + the additional 20 minutes = 4:20pm

OpenStudy (anonymous):

thank you!

OpenStudy (gw2011):

You're welcome